OpenStudy (clara1223):

find the limit as x approaches 0 of (tan(3x))/x (Hint: write tangent in terms of sine and cosine) a) 1/3 b) 3 c) 0 d) 1 e) does not exist

zepdrix (zepdrix):

Sines and cosines? :o Oh ok, this will be almost identical to the last problem then.

zepdrix (zepdrix):

\[\large\rm \lim_{x\to 0}\frac{\tan(3x)}{x}=\lim_{x\to 0}\frac{\frac{\sin(3x)}{\cos(3x)}}{x}=\lim_{x\to 0}\frac{\sin(3x)}{x \cos(3x)}\]Going further,\[\large\rm =\lim_{x\to 0}\frac{\sin(3x)}{x}\cdot\lim_{x\to 0}\frac{1}{\cos(3x)}\]

zepdrix (zepdrix):

You ok with that third equals sign? I know that little algebra step can be tricky.

OpenStudy (clara1223):

yes, I understand

OpenStudy (clara1223):

so the answer again is just 3 times 1?

zepdrix (zepdrix):

yay \c:/